OpenStudy (anonymous):

Write an equation for the line that passes through the points (5, -1) and (-7, -8). A. y = 7/12 x – 47/12 B. y = 7/12 x + 67/12 C. y = 12/7 x – 67/7 D. y = 12/7 x – 4

OpenStudy (luigi0210):

Find the slope first: \[\LARGE m=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 5}}\quad ,&{\color{blue}{ -1}})\quad &({\color{red}{ -7}}\quad ,&{\color{blue}{ -8}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)

OpenStudy (anonymous):

i got the -7/-12

OpenStudy (luigi0210):

Use that and plug it into the point-slope formula: \[\LARGE y-y_{1}=m(x-x_{1})\] jdoe did the points :P

OpenStudy (triciaal):

one approach find the slope of the line change in y / change in x using one of the given points find the intercept b using y = mx + b you have m and b write the equation

OpenStudy (anonymous):

okay sorry i was taking forever i am being a fat *** and eating chips

OpenStudy (anonymous):

\[-7--1=m \left( -12-5 \right)\]

OpenStudy (triciaal):

did you get a

OpenStudy (anonymous):

idk

OpenStudy (triciaal):

what is the slope?

OpenStudy (anonymous):

\[\frac{ -7 }{ -12 }\]